### v1

# JWST/MIRI-MRS view of the metal-poor galaxy CGCG 007-025: the spatial location of PAHs and very highly ionized gas

Abstract
Polycyclic Aromatic Hydrocarbons (PAHs) are key diagnostics of the physical conditions in the interstellar medium and are widely used to trace star formation in the mid-infrared (mid-IR). The relative strengths of mid-IR PAH emission features (e.g., 6.2, 7.7, 11.3 m) are sensitive to both the size and ionization state of the molecules and can be strongly influenced by the local radiation field. However, at low metallicities (), detecting PAHs remains notoriously difficult, likely reflecting a combination of suppressed formation and enhanced destruction mechanisms. We present new JWST/MIRI MRS observations of the metal-poor () dwarf galaxy CGCG 007-025. We confirm the tentative PAH detection previously reported from Spitzer data and, for the first time, identify a compact ( pc) PAH-emitting region nearly co-spatial with the newly detected [Ne V] (I.P. eV) emission and the galaxy’s most metal-poor, strongly star-forming region. The feature is clearly detected, while no emission is found from the other typically brighter features, suggesting a PAH population dominated by large, neutral molecules resilient to hard ionizing fields. When compared with models, mid-IR line ratios involving [Ne III], [O IV], and [Ne V] can only be reproduced by a combination of star formation and AGN ionization, with the latter contributing 4–8%. The [O IV] and [Ne V] luminosities exceed what massive stars or shocks can produce, highlighting a puzzling scenario in line with recent JWST observations of similar galaxies. This work provides a crucial reference for studying the physical conditions of the dust and star formation in low-metallicity starburst regions, environments typical of the early universe.
1 Introduction
Polycyclic Aromatic Hydrocarbons (PAHs) are widespread components of the interstellar medium (ISM), producing strong vibrational features in the near- and mid-infrared (IR) at 3.3, 6.2, 7.7, and 11.3 (see Li, 2020, for a review). These features often dominate the mid-IR spectra of galaxies and serve as key tracers of the physical conditions in the ISM and star formation intensity (Sales et al., 2010; Shipley et al., 2016, among others). More importantly, PAHs have been used as tracers of the dust content and composition in galaxies (e.g., Haas et al., 2002; Jones et al., 2015; Sutter et al., 2024). Understanding their role is particularly crucial in low-metallicity environments ( Wu et al., 2006; Hunt et al., 2010), where the origin and survival of dust remain poorly constrained. A recent JWST study by Shivaei et al. (2024) has found that the PAH mass fraction in galaxies at –2 declines sharply below and , consistent with a combination of enhanced PAH destruction in strong radiation fields and reduced efficiency of PAH formation at low metallicities. This picture has been challenged by recent observations of high-redshift galaxies. Since the metallicity of the overall galaxy population decreases with redshift (e.g., Curti et al., 2023), one might expect low dust content in the early Universe. However, observations of unexpectedly dusty galaxies at very high redshift (up to ) indicate that substantial dust reservoirs were already in place when asymptotic giant branch stars could not yet have contributed significantly to the dust budget (Witstok et al., 2023; Ormerod et al., 2025), suggesting that both the production and destruction channels of dust and complex molecules such as PAHs are not yet fully understood.
In this letter, we focus on the nearby (; , Kourkchi et al., 2020), metal deficient (% Z⊙, Izotov & Thuan, 2004; van Zee & Haynes, 2006; Izotov et al., 2007, among others) dwarf galaxy (, Marasco et al., 2023) CGCG 007-025. This starbursting galaxy displays a prominent star-forming region with and [O III]5007/[O II]3727 2.9, satisfying the criteria to be treated as a local analogue of high-redshift galaxies (e.g., Berg et al., 2022; Brammer et al., 2012). Previous observations of this galaxy with Spitzer/IRS revealed the presence of very weak PAH features located at 7.7 and 11.3 (Hunt et al., 2010). Using new JWST MIRI/MRS data, we uniquely characterize the properties of PAH emission features in CGCG 007-025 to provide new insights into the formation and destruction mechanisms of these molecules in low-metallicity environments, similar to those of the first galaxies in the Universe.
In Section 2 we describe the MIRI/MRS observations and data reduction, and we present our emission line measurements and PAH modeling methodology in Section 3. The main results on the spatial extension of the PAH features, together with the general properties of the emitting region, are compiled in Section 4. The discussion is in Section 5. We summarize our findings in Section 6.
2 Data
This letter presents the JWST MIRI/MRS data of CGCG 007-025. The main starbursting region of the galaxy was observed as part of the Cycle 2 JWST program GO 4278 (PI: Mingozzi). Figure 1 shows the MIRI/MRS field of view (FoV) coverage superimposed on the Subaru/HSC color image. The data were reduced using the JWST data reduction pipeline version 1.17.1.dev with the CRDS 1338 context (see Appendix A for more details). The final data products were the 12 sub-channel cubes with the default spatial and spectral sampling. The data show a PSF-like structure (see inset of Figure 1), which causes fringe patterns to vary systematically across the PSF and requires further corrections (see Mingozzi et al., 2025, for a full description of the mitigation process we applied).
3 Methodology
Depending on the signal-to-noise ratio (S/N) of the continuum, we adopted two complementary approaches to analyze the mid-IR spectral features of CGCG 007-025. For the integrated spectrum, where the S/N is high (S/N in the continuum), we modeled the continuum and spectral features using pahfit (Smith et al., 2007). At the spaxel level, where the continuum S/N is lower, we instead performed localized fits in narrow spectral windows around each feature of interest.
3.1 Extended ionized gas and PAH emission
To characterize the extended emission of the ionized gas at the spaxel level, we used an approach similar to the one described in Mingozzi et al. (2025). We used a combination of a Gaussian profile and a 1-degree polynomial –to account for emission-line and surrounding continuum emission– within a km/s window around each line of interest. We fit the emission lines listed in Table 1.
To characterize the extension of the PAH emission, we focus our analysis on the 11.3 bump –the brightest and only visible PAH feature (see Section 4.2). We created a sliced cube (from the original channel 2 cube) of km/s around this feature and applied a moving median smooth filter over 71 spectral pixels to reduce the noise on the bump (similarly to Hunt et al., 2025), since this feature appears especially weak when using individual spaxels. After smoothing, we modeled the spectra using a combination of a 1-degree polynomial and a Drude profile (the standardized profile for PAH emission, see Draine & Li, 2007).
3.2 Integrated emission
Prior to extraction, we matched the spatial resolution across bands using the linear relation between the MIRI/MRS PSF FWHM and wavelength from Law et al. (2023) as a first approximation for the PSF variation between channels. At each wavelength step, we convolved the data with a 2D Gaussian kernel whose width matches the difference in quadrature between the PSF FWHM at that wavelength and the reference value. We considered the reference value to be that of [O IV] 25.89, corresponding to a px (095, 98 pc). After convolution, we resampled the cubes to match the pixel scale of channel 1 (013).
Once the integrated spectrum is on hand, we use pahfit to model the mid-IR continuum of CGCG 007-025. pahfit fits simultaneously the dust continuum, PAH features and emission lines using a combination of blackbody, Drude and Gaussian profiles, respectively, together with an extinction curve (Smith et al., 2007). We update the default list of emission lines to be modeled by pahfit using the line list presented in Van De Putte et al. (2024). Moreover, we allowed the PAH profiles to vary in central wavelength (1%) and in FWHM (-60%, +10%), similarly to Donnan et al. (2023).
4 Results
4.1 Location and extension of the 11.3 PAH feature
We present, in the left panel of Figure 2, the spatially resolved 11.3 PAH feature. The 11.3 PAH emission peaks near the location of one of the clumps in CGCG 007-025, with an extension of 50 pc (delimited by the pink contour). At this wavelength, the native PSF has a = 2.8 px (048, 48 pc), so that the region is barely spatially resolved in channel 2. We used this contour to extract a spectrum which maximizes the signal of the 11.3 feature (see right panel of Figure 2).
4.2 Integrated spectrum of the PAH emitting region
Following the method described in Section 3.2, and using the aperture identified on the PAH map, we extracted the spectrum of the PAH emitting region (solid gray line in the right panel of Figure 2). The best fit model from pahfit, together with the residuals, are displayed in the same figure in light orange and light green, respectively. We also zoom around the 11.3 PAH to highlight the performance of pahfit in reproducing this feature. In Table 2 (see Appendix B) we compile all the line fluxes measured by pahfit. These fluxes have been corrected for dust attenuation by , where is the attenuation fraction from the pahfit model (dashed gray line in right panel of Figure 2)
From the integrated spectrum shown in Figure 2, there is no clear evidence of any other PAH emission in CGCG 007-025. However, PAH features such as the 6.2, 7.7 or 8.6 are typically among the brightest ones (e.g., Hunt et al., 2010; Sandstrom et al., 2012; Chastenet et al., 2019). Following García-Bernete et al. (2022b), we obtain the upper limit on the 6.2 feature using the residuals from the pahfit model over the wavelength range 6.0 to 6.5 . We also report the upper limit on the 6.2 feature in Table 2.
4.3 Ionized gas emission in the mid-IR
The integrated spectrum of the PAH-emitting region presented in Figure 2 reveals the detection of a plethora of nebular emission lines such as [Fe II] 5.34, [Ar II] 6.99, [Ne II] 12.81, [S III] 18.71, [Ar III] 8.99, [S IV] 10.51, [Ne III] 15.55, and [O IV] 25.89, as well as the high ionization line [Ne V]111Here we focus only on [Ne V] 14.32, since the accompanying line at 24.32 falls at the edge of ch4-medium and ch4-long, challenging the detection of the feature. 14.32. Accordingly, we search for the detection of other lines at the spaxel level, including lines of high ionization potential (IP) such as [O IV] and [Ne V]. More concretely, following the methods outlined in Section 3.1, we focus on [Ne II], [Ne III], [O IV] and [Ne V] due to the diagnostic power of these lines (see Section 5). While [Ne II] and [Ne III] emission extend across the entire MIRI FoV (see inset in Figure 1), both [O IV] and [Ne V] emission are in a more confined configuration, co-spatial with the same compact region as the extended emission (see light green contour in Figure 2).
4.4 Previous observations with Spitzer
As part of an analysis on Blue Compact Dwarfs (BCDs) with Spitzer/IRS, Hunt et al. (2010) studied the infrared spectrum of CGCG 007-025. Combining the IRS low and high resolution gratings, they analyzed the integrated spectrum of the brightest star-forming region, and reported the detection of two PAH features: the 7.7 and 11.3 complexes. They also reported the detection of the fine structure lines [Ar III], [S IV], [Ne III] and [S III], together with some lines.
We extracted the integrated MIRI/MRS spectrum of CGCG 007-025 using the methodology outlined in Section 3.2, with the spatial footprint of MIRI/MRS channel 1 (FoV 32 × 37) as the defining aperture to try to match the Spitzer/IRS aperture (see Figure 1). The top panel of Figure 3 displays this spectrum, with the different emission lines identified. In order to visually compare it with the literature, we degraded the MIRI/MRS data to match the spectral resolution of the IRS spectrum from Hunt et al. (2010). We convolved the MIRI/MRS integrated spectrum with the Spitzer/IRS LSF, and then binned the spectrum to a 0.03 wavelength step. This spectrum is shown as a maroon thick-line in both panels of Figure 3. The faint emission lines in the integrated MIRI/MRS spectrum (such as [Ar II] or [Ne V]) are not visible at the Spitzer sensitivity. Moreover, lines that are close to each other are blended at the Spitzer resolution. This is the case of the Pfund () and Humphrey () lines, which conspire to show up as a small bump at the position of the 7.7 PAH feature, clearly absent in MIRI/MRS high-resolution spectrum. This highlights the superior capability of MIRI/MRS compared to Spitzer to resolve mid-IR individual features.
In the bottom panel of Figure 3 we show together the degraded MIRI/MRS spectrum with the reference spectrum from Hunt et al. (2010). We note a discrepancy in the spectral slopes, which may be due to a slight aperture mismatch and/or the specific sensitivity characteristics of MIRI. We also emphasized in green the theoretical central wavelengths of the PAH features. The two PAH complexes reported in Hunt et al. (2010) are marked in dark green. Based on this visual comparison and given the high spectral resolution and sensitivity of MIRI/MRS, we argue that only the 11.3 m feature is present in the galaxy, while the 7.7 m feature is not detected within the covered FoV.
5 Discussion
5.1 PAHs as star formation rate indicators
PAH emission has long been proposed to be a potential star formation rate (SFR) indicator. Using Spitzer data, Xie & Ho (2019) presented a relation between the PAH luminosity and the SFR obtained from [Ne II] and [Ne III] for different types of galaxies, including BCDs, star-forming galaxies (SFGs) and AGN (active galactic nuclei). Although the PAH intensity is suppressed in the low-mass () galaxies, the Xie & Ho (2019) SFR-PAH calibrations can still be applied to this mass regime. Our MIRI/MRS data allows us to test for the validity of PAH features as SFR tracers, not only at low metallicity but also in a spatially resolved way.
In order to estimate the SFR in the PAH emitting region, we used the prescription presented in Zhuang et al. (2019), which is an update of Ho & Keto (2007) to include the effects of metallicity. This estimator relies on the [Ne II] and [Ne III] luminosities and their fractional abundances, and includes a correction to account for the presence of [Ne V]. The use of [Ne II] and [Ne III] to calculate the SFR is motivated by the lack of a reliable estimator based solely on mid-IR hydrogen lines. Moreover, this SFR tracer has been calibrated against the Paschen- emission, making it more suitable for the case explored here. Using the extinction corrected values from pahfit and a metallicity of (Berg et al., 2022; Fernández et al., 2023), we obtain a in the PAH emitting region. When compared with the literature, the point in CGCG 007-025 falls within the scatter of the other BCDs from Xie & Ho (2019) (see Figure 6 in Appendix B), and within the relation between for BCDs. However, this measurement only encompasses the region in which the PAH emission is reliably measured while, intriguingly, the area enclosing the SFR peak (solid white contour in Figure 2) does not display significant PAH emission. It remains possible that these calibrations do not hold in low-metallicity environments when using spatially resolved data.
5.2 What PAH and line ratios reveal about CGCG 007-025
PAH features at different wavelengths can be used to trace the predominant ionization state of the overall PAH population. The emission features at 6.2 and 7.7 are typically associated with ionized PAHs, while the feature at 11.3 is emitted by large, neutral PAHs. Consequently, the ratios between PAH features provide insight into the dominant PAH population and the ionizing mechanisms within a galaxy. For instance, several studies have reported thresholds in the 7.7/11.3 and 6.2/11.3 ratios, with SF-dominated regions typically showing 6.2/11.3 larger than 1, while environments exposed to harder radiation fields (e.g., AGN) display ratios below 0.3 (e.g., Ogle et al., 2010; García-Bernete et al., 2022a; Donnan et al., 2023).
We show in Figure 4 the 6.2/11.3 PAH ratio against [Ne III]/[Ne II], which is tracing the level of ionization. Local HII galaxies are concentrated in the upper left corner of the plot, with LIRGs (Luminous Infrared Galaxies) displaying similar [Ne III]/[Ne II] ratios but slightly lower 6.2/11.3 ratios. Seyfert galaxies move towards the center of the plot, showing an increase in the level of ionization and a further decrease in the 6.2/11.3 ratio. BCDs, on the other hand, cover a wider PAH ratio range while showing even higher ionization. CGCG 007-025 (orange arrow in the lower right corner) stands out as an outlier in this diagram, showing one of the largest [Ne III]/[Ne II] ratio together with a deficit on the 6.2/11.3 PAH ratio. Other BCDs, like IIZw40 (; Lai et al. 2025), display a larger [Ne III]/[Ne II] ratio, but their PAH ratios, and metallicity, are not as low as in the case of CGCG 007-025. The location of CGCG 007-025 in the diagram may indicate the presence of a highly ionizing source, as its PAH strengths are consistent with those typically observed in AGN-dominated systems. Indeed, its PAH and mid-IR line properties closely resemble those observed in the nucleus of the Seyfert 2 galaxy NGC 7319 and in radio galaxies with strong jets, which show only a marginal detection of the 11.3 m PAH feature (Pereira-Santaella et al., 2022; García-Bernete et al., 2022b; Zakamska et al., 2016). Similarly, the [Ne III]/[Ne II] and PAH 6.2/11.3 ratios of the AGN NGC 5728 (García-Bernete et al., 2024) are consistent with those measured in this galaxy.
In addition to a significantly large [Ne III]/[Ne II] ratio, the integrated spectrum of CGCG 007-025 also displays high-ionization mid-IR lines such as [O IV] (I.P. 50 eV) and [Ne V] (I.P. 97 eV), which are hard to explain in pure star-forming environments (e.g., Lutz et al., 1998; Schaerer & Stasińska, 1999). Nevertheless, [O IV] emission is commonly observed in BCDs (e.g., Dale et al., 2009; Bernard-Salas et al., 2009), and recent MIRI/MRS observations have revealed extended [Ne V] in some extreme low-metallicity systems, such as SBS 0335–052E and I Zw 18 (Mingozzi et al. 2025; Hunt et al. 2025; see also e.g., Izotov et al. 2012, 2021 for detections on the optical). The origin of such high-ionization emission in low-metallicity environments remains puzzling and not yet fully understood (Berg et al., 2021; Olivier et al., 2022). We further explore this in Figure 5, which displays the [Ne V]/[Ne II] vs [Ne III]/[Ne II] (top panel) and [O IV]/[Ne III] vs [Ne III]/[Ne II] (bottom panel) excitation diagnostic diagrams. CGCG 007-025 PAH emitting region and emission line ratios at spaxel level are shown as a star and colored dot symbols, respectively, and color-coded in accordance to the complementary line ratio of each panel (the value of the star is marked in the colorbar). Overlaid in green are the single stellar population (SSP) models from Martínez-Paredes et al. (2023) for an age222This is the only age grid that appears within the displayed [Ne V]/[Ne II], [Ne III]/[Ne II] and [O IV]/[Ne III] ranges and is the closest to the estimated age for CGCG 007-025 of 5 Myr from del Valle-Espinosa et al. (2023) of 2.5 Myr, with a top-heavy IMF and and metallicity ranging from . Interestingly, in both panels, the data points are far away from the locus of the pure SF models.
For the other two recently discovered [Ne V]-emitters - SBS 0335–052E (light blue markers and contours in Figure 5) and I Zw 18 - the only models capable of reproducing the observed line ratios require the presence of an additional source, such as an intermediate-mass black hole (IMBH; Mingozzi et al., 2025; Hunt et al., 2025). Thus, we included in Figure 5 the IMBH grids from Richardson et al. (2022) with and the youngest age available (20 Myr), color-coded by the AGN fraction. Consistently with Mingozzi et al. (2025), the line ratios of CGCG 007-025 could be explained with a modest 4-8% of AGN fraction. Another possible source of ionization capable of generating [Ne V] could be shocks. We have included in Figure 5, in black, the shock+precursor models from Flury et al. (2025) with 10% . However, these models fail to reproduce, simultaneously, the high and low ionization species seen in CGCG 007-025. Overall, we would like to remark this is the first time that the emission line ratios of this galaxy fall outside the SF locus in diagnostic diagrams (using e.g., UV and optical line ratios; Mingozzi et al., 2022; del Valle-Espinosa et al., 2023). It remains possible that current stellar population models underestimate the production of high-energy ionizing photons, particularly in low-metallicity environments (e.g., Sellmaier et al., 1996; Stasińska et al., 2015; Fernández-Ontiveros et al., 2016). The present dataset therefore offers a valuable benchmark for guiding and constraining the next generation of star-formation models aimed at capturing such extreme conditions.
Complementing these diagnostics, García-Bernete et al. (2024) recently studied the relation between the PAH feature at 6.2 and the S(1) line to establish alternative ways to determine the presence of an AGN in the absence of high IP lines. The authors suggest that ratios of 6.2/S(1) larger than 12 are dominated by SF processes, whereas ratios below this value reflect an increase on the due to the heating of shocks/AGN. Stronger radiation fields would further destroy the PAHs, reducing the ratio below 5. For CGCG 007-025, we report a lower limit of , which would fall in the AGN demarcation. Nevertheless, this trend has been established in (super-)solar metallicity environments as well as in confirmed AGN host galaxies.At lower metallicities, it is still to be determined whether this ratio can be attributed to the harder ionizing spectrum (which would destroy the PAH complexes and increase the S(1) flux, simultaneously), or whether they are shaped simply because of the lack of PAH molecules in the first place.
It is worth noting that the PAH emission peaks at the location of high-IP lines ([O IV] and [Ne V]), with no significant PAH detections in regions outside of these high-IP boundaries. In the scenario of PAH destruction in the presence of a highly ionizing field, this spatial anti-correlation seems contradictory: smaller PAHs should survive preferentially in lower-ionization environments. An alternative to PAH destruction is in-situ formation, where the source producing [O IV] and [Ne V]emission could also be the origin of PAH molecules. For example, the formation of PAH has previously been seen in the dusty winds of massive and/or Wolf-Rayet (WR) stars (Lau et al., 2022; Taniguchi et al., 2025). Although WR stars seem capable of producing [Ne V] emission (see Tarantino et al., 2024; Hunt et al., 2025), current photoionization models of massive stars cannot reproduce the mid-IR line ratios seen in CGCG 007-025. Further theoretical and observational efforts are thus required to understand the production of high-ionization photons and the formation of dust and complex molecules in low-metallicity starburst environments.
6 Conclusions
In this letter, we present the first clear and spatially resolved PAH detection at in the dwarf galaxy CGCG 007-025 using the unique capabilities of JWST MIRI/MRS. Our findings are summarized below:
The 11.3 PAH feature is significantly detected at spaxel level (left panel of Figure 2). We explore the extension and location of this feature. The emission is spatially confined to a region of pc in the vicinity of the brightest clumps on the galaxy.
Using pahfit, we characterize the continuum and spectral features (nebular lines and PAHs) of the integrated spectrum of the PAH emitting region (right panel of Figure 2). We confirm the presence of the 11.3 PAH feature with a flux of .
We do not find evidence of any other PAH feature in the integrated spectrum of CGCG 007-025. We place an upper limit detection on the 6.2 PAH complex of .
The exclusive detection of the 11.3 feature indicates a PAH population dominated by larger, neutral molecules. Moreover, the 6.2/11.3 ratio, in conjunction with a large [Ne III]/[Ne II] ratio (Figure 4), hint to the presence of a strong ionizing source, with [Ne III]/[Ne II] ratios compatible with those seen in harsh environments like AGNs or other BCDs.
The presence of [Ne V] and [O IV] in the region is hard to explain using pure SF models. Instead, the line ratios are compatible with a mix of SF+AGN photoionization, with a moderate AGN fraction no larger than 8% (Figure 5). However, it remains possible that stellar models fail to produce enough high-ionizing photons.
To conclude, this study demonstrates that JWST/MIRI data is capable of not only detecting but also spatially resolving faint PAH emission in metal-poor environments. As a result, this instrument provides new insight into the effectiveness of mid-IR diagnostics and the nature of ionizing sources in systems previously considered to be purely star-forming systems, offering valuable guidance for understanding the physical conditions of dust and gas in metal-poor environments typical of the high-z universe.
Appendix A Data reduction
We retrieved the mid-IR data of CGCG 007-025 from MAST. The galaxy was observed with JWST/MIRI MRS using the three MRS grating settings (SHORT, MEDIUM and LONG) and the FAST readout mode, giving full coverage in the wavelength range . A single pointing was taken, centered in the main starbursting region of the galaxy. Given the JWST pointing accuracy of ”, target acquisition was not required. A 4-point dither pattern was implemented to ensure optimal sampling across the FoVs and to facilitate the identification and removal of detector artifacts. Each pointing was observed with 50 groups per integration, with a total exposure time of 1443 seconds. To enable accurate background subtraction, dedicated background observations were conducted using half the number of dithers. The background and science observations were placed in an uninterruptible sequence.
The data were reduced using the JWST data reduction pipeline version 1.17.1.dev with the CRDS 1338 context. Briefly, in Stage 1 corrections at detector level are applied to produce data in count rate units. These include corrections for dark current, masking of detector cosmetics, and removal of strong jumps introduced by the impact of cosmic rays. In Stage 2, data is flat-field corrected, and wavelength and flux calibrated. A first inspection of the count rate images revealed intense cosmic ray events all across the detector which are difficult to mask. At the time of reducing the data, the developer version of the pipeline was the only release which included the straylight correction step, a crucial step for correcting these intense cosmic ray events. Background subtraction is performed in Stage 3 using a master sky background created with the dedicated sky observations of each object. We opted for this option rather than a pixel-by-pixel correction, since the later introduces more noise in the final data products. Stage 3 also combines the individual exposures into a single one and reconstructs the 3D datacubes per band. We also enabled a residual fringe correction in Stage 2 to account for any residual fringing not corrected during Stage 1. For extended sources, the simple fringe-flat correction applied in Stage 1 is usually enough to reduce its impact on the data products. However, the galaxy displays a PSF-like structure, due to an unresolved compact emission, making the fringe patterns vary systematically across the PSF and requiring further corrections to create the final extended mid-IR emission datacubes. We chose the emsm algorithm as a first step to minimize the wiggle patterns of the PSF. The additional PSF-like emission modeling and subtraction was done following the Mingozzi et al. (2025) procedure. Besides our best efforts to correct from fringing and PSF undersampling, some spaxels situated at the peak of the mid-IR emission are still affected by these wiggling patterns.
Appendix B Additional material
In Table 1 we compile the complete list of emission lines included in pahfit for the modeling of the PAH emitting region. Table 2 displays the unattenuated emission line measurements from pahfit together with the flux of the 11.3 m PAH feature and the upper limit on the 6.2 m PAH feature.
| Fine structure lines | H I lines | lines | ||||
| Line name | IP | Transition | Line name | |||
| [eV] | [] | [] | [] | |||
| [Fe II] | 7.9 | 5.34 | Hu(10-6) | 5.129 | S(8) | 5.053 |
| [Ar II] | 15.76 | 6.99 | Hu(9-6) | 5.908 | S(7) | 5.511 |
| [Ar III] | 27.63 | 8.99 | Pf(6-5) | 7.45985 | S(6) | 6.109 |
| [S IV] | 34.83 | 10.51 | Hu(8-6) | 7.503 | S(5) | 6.909 |
| [Ne II]⋆ | 21.56 | 12.81 | 10-7 | 8.76 | S(4) | 8.026 |
| [Ne V]⋆ | 97.12 | 14.32 | 9-7 | 11.3088 | S(3) | 9.665 |
| [Ne III]⋆ | 40.96 | 15.55 | Hu(7-6) | 12.37188 | S(2) | 12.278 |
| [S III] | 23.34 | 18.71 | 8-7 | 19.062 | S(1) | 17.035 |
| [O IV]⋆ | 54.95 | 25.89 | ||||
| [Fe II] | 7.9 | 25.99 | ||||
| Fine structure lines | lines | PAH | |||
| Line name | Flux | Line name | Flux | Feature | Flux |
| [Fe II]5.34 | 0.4 0.3 | S(1) | 0.80 0.03 | 6.2 | |
| [Ar II]6.99 | 0.29 0.08 | 11.3 | 5.1 0.4 | ||
| [Ar III]8.99 | 3.86 0.08 | ||||
| [S IV]10.51 | 31.2 0.4 | ||||
| [Ne II]12.81 | 2.22 0.03 | ||||
| [Ne V]14.32 | 0.38 0.02 | ||||
| [Ne III]15.55 | 35.3 0.4 | ||||
| [S III]18.71 | 43.4 0.4 | ||||
| [O IV]25.89 | 24.4 0.4 | ||||
| [Fe II]25.99 | 1.2 0.3 | ||||
In Figure 6 we show the SFR-PAH relation for the 11.3 (left) and 6.2 (right) PAH features. The BCDs sample from Xie & Ho (2019) is displayed as purple diamonds, with open symbols representing upper limits on the PAH luminosities. The different SFR-PAH calibrations from Xie & Ho (2019) are shown in purple, with other calibrations from the literature displayed in green (Mordini et al., 2021). The measurements from CGCG 007-025 are shown in orange.

### v2

@@
 Polycyclic Aromatic Hydrocarbons (PAHs) are key diagnostics of the physical conditions in the interstellar medium and are widely used to trace star formation in the mid-infrared (mid-IR).
-The relative strengths of mid-IR PAH emission features (e.g., 6.2, 7.7, 11.3 m) are sensitive to both the size and ionization state of the molecules and can be strongly influenced by the local radiation field.
-However, at low metallicities (), detecting PAHs remains notoriously difficult, likely reflecting a combination of suppressed formation and enhanced destruction mechanisms.
-We present new JWST/MIRI MRS observations of the metal-poor () dwarf galaxy CGCG 007-025.
+The relative strengths of mid-IR PAH features (e.g., 6.2, 7.7, 11.3 m) are sensitive to both the size and ionization state of the molecules and can be strongly influenced by the local radiation field.
+However, at low metallicities (Z0.2Z⊙), detecting PAHs remains notoriously difficult, likely reflecting a combination of suppressed formation and enhanced destruction mechanisms.
+We present new JWST/MIRI MRS observations of the metal-poor (Z0.1Z⊙) dwarf galaxy CGCG 007-025.
 We confirm the tentative PAH detection previously reported from Spitzer data and, for the first time, identify a compact ( pc) PAH-emitting region nearly co-spatial with the newly detected [Ne V] (I.P.
 eV) emission and the galaxy’s most metal-poor, strongly star-forming region.
-The feature is clearly detected, while no emission is found from the other typically brighter features, suggesting a PAH population dominated by large, neutral molecules resilient to hard ionizing fields.
+The PAH11.3μm and PAH12.7μm features are detected, while no emission is found from other typically brighter features, suggesting a PAH population dominated by large, neutral molecules resilient to hard ionizing fields.
 When compared with models, mid-IR line ratios involving [Ne III], [O IV], and [Ne V] can only be reproduced by a combination of star formation and AGN ionization, with the latter contributing 4–8%.
@@
 This work provides a crucial reference for studying the physical conditions of the dust and star formation in low-metallicity starburst regions, environments typical of the early universe.
-1 Introduction
+I Introduction
 Polycyclic Aromatic Hydrocarbons (PAHs) are widespread components of the interstellar medium (ISM), producing strong vibrational features in the near- and mid-infrared (IR) at 3.3, 6.2, 7.7, and 11.3 (see Li, 2020, for a review).
@@
 More importantly, PAHs have been used as tracers of the dust content and composition in galaxies (e.g., Haas et al., 2002; Jones et al., 2015; Sutter et al., 2024).
-Understanding their role is particularly crucial in low-metallicity environments ( Wu et al., 2006; Hunt et al., 2010), where the origin and survival of dust remain poorly constrained.
+Understanding their role is particularly crucial in low-metallicity environments (, Wu et al., 2006; Hunt et al., 2010), where the origin and survival of dust remain poorly constrained.
 A recent JWST study by Shivaei et al.
@@
 However, observations of unexpectedly dusty galaxies at very high redshift (up to ) indicate that substantial dust reservoirs were already in place when asymptotic giant branch stars could not yet have contributed significantly to the dust budget (Witstok et al., 2023; Ormerod et al., 2025), suggesting that both the production and destruction channels of dust and complex molecules such as PAHs are not yet fully understood.
-In this letter, we focus on the nearby (; , Kourkchi et al., 2020), metal deficient (% Z⊙, Izotov & Thuan, 2004; van Zee & Haynes, 2006; Izotov et al., 2007, among others) dwarf galaxy (, Marasco et al., 2023) CGCG 007-025.
+In this letter, we focus on the nearby (; , Kourkchi et al., 2020), metal deficient (% Z⊙ or 12 + log(O/H) 7.7, Izotov and Thuan, 2004; van Zee and Haynes, 2006; Izotov et al., 2007, among others) dwarf galaxy (, Marasco et al., 2023) CGCG 007-025.
 This starbursting galaxy displays a prominent star-forming region with and [O III]5007/[O II]3727 2.9, satisfying the criteria to be treated as a local analogue of high-redshift galaxies (e.g., Berg et al., 2022; Brammer et al., 2012).
 Previous observations of this galaxy with Spitzer/IRS revealed the presence of very weak PAH features located at 7.7 and 11.3 (Hunt et al., 2010).
-Using new JWST MIRI/MRS data, we uniquely characterize the properties of PAH emission features in CGCG 007-025 to provide new insights into the formation and destruction mechanisms of these molecules in low-metallicity environments, similar to those of the first galaxies in the Universe.
-In Section 2 we describe the MIRI/MRS observations and data reduction, and we present our emission line measurements and PAH modeling methodology in Section 3.
-The main results on the spatial extension of the PAH features, together with the general properties of the emitting region, are compiled in Section 4.
-The discussion is in Section 5.
-We summarize our findings in Section 6.
-2 Data
+Using new JWST MIRI/MRS data, we uniquely characterize the properties of PAH emission features in CGCG 007-025 to provide new insights into the survivability of these molecules in low-metallicity environments, similar to those of the first galaxies in the Universe.
+In Section II we describe the MIRI/MRS observations and data reduction, and we present our emission line measurements and PAH modeling methodology in Section III.
+The main results on the spatial extension of the PAH features, together with the general properties of the emitting region, are compiled in Section IV.
+The discussion is in Section V.
+We summarize our findings in Section VI.
+II Data
 This letter presents the JWST MIRI/MRS data of CGCG 007-025.
@@
 The data were reduced using the JWST data reduction pipeline version 1.17.1.dev with the CRDS 1338 context (see Appendix A for more details).
-The final data products were the 12 sub-channel cubes with the default spatial and spectral sampling.
+The final data products were the 12 sub-channel cubes with the default spatial and spectral sampling, whose astrometry matches with HST/WFC3 F814W filter.
 The data show a PSF-like structure (see inset of Figure 1), which causes fringe patterns to vary systematically across the PSF and requires further corrections (see Mingozzi et al., 2025, for a full description of the mitigation process we applied).
-3 Methodology
+III Methodology
 Depending on the signal-to-noise ratio (S/N) of the continuum, we adopted two complementary approaches to analyze the mid-IR spectral features of CGCG 007-025.
@@
 At the spaxel level, where the continuum S/N is lower, we instead performed localized fits in narrow spectral windows around each feature of interest.
-3.1 Extended ionized gas and PAH emission
+III.1 Extended ionized gas and PAH emission
 To characterize the extended emission of the ionized gas at the spaxel level, we used an approach similar to the one described in Mingozzi et al.
@@
 We fit the emission lines listed in Table 1.
-To characterize the extension of the PAH emission, we focus our analysis on the 11.3 bump –the brightest and only visible PAH feature (see Section 4.2).
-We created a sliced cube (from the original channel 2 cube) of km/s around this feature and applied a moving median smooth filter over 71 spectral pixels to reduce the noise on the bump (similarly to Hunt et al., 2025), since this feature appears especially weak when using individual spaxels.
-After smoothing, we modeled the spectra using a combination of a 1-degree polynomial and a Drude profile (the standardized profile for PAH emission, see Draine & Li, 2007).
-3.2 Integrated emission
+To characterize the extension of the PAH emission, we focus our analysis on the 11.3 bump –the brightest PAH feature present in this object (see Section IV.2).
+We created a sliced cube (from the original channel 2 cube) of km/s around this feature and applied a moving median smooth filter over 71 spectral pixels to reduce the noise on the bump (similarly to Hunt et al., 2025b), since this feature appears especially weak when using individual spaxels.
+After smoothing, we modeled the spectra using a combination of a 1-degree polynomial and a Drude profile (the standardized profile for PAH emission, see Draine and Li, 2007).
+III.2 Integrated emission
 Prior to extraction, we matched the spatial resolution across bands using the linear relation between the MIRI/MRS PSF FWHM and wavelength from Law et al.
@@
 (2023).
-4 Results
-4.1 Location and extension of the 11.3 PAH feature
+IV Results
+IV.1 Location and extension of the 11.3 PAH feature
 We present, in the left panel of Figure 2, the spatially resolved 11.3 PAH feature.
@@
 At this wavelength, the native PSF has a = 2.8 px (048, 48 pc), so that the region is barely spatially resolved in channel 2.
+We also show, in pink, the flux distribution of this feature in the bottom left panel of Figure 2.
 We used this contour to extract a spectrum which maximizes the signal of the 11.3 feature (see right panel of Figure 2).
-4.2 Integrated spectrum of the PAH emitting region
-Following the method described in Section 3.2, and using the aperture identified on the PAH map, we extracted the spectrum of the PAH emitting region (solid gray line in the right panel of Figure 2).
-The best fit model from pahfit, together with the residuals, are displayed in the same figure in light orange and light green, respectively.
+IV.2 Integrated spectrum of the PAH emitting region
+Following the method described in Section III.2, and using the aperture identified on the PAH map, we extracted the spectrum of the PAH emitting region (solid gray line in the right panel of Figure 2).
+The best fit model from pahfit, together with the residuals, are displayed in the same figure in light orange and light purple, respectively.
 We also zoom around the 11.3 PAH to highlight the performance of pahfit in reproducing this feature.
-In Table 2 (see Appendix B) we compile all the line fluxes measured by pahfit.
-These fluxes have been corrected for dust attenuation by , where is the attenuation fraction from the pahfit model (dashed gray line in right panel of Figure 2)
+In Table 1 (see Appendix C) we compile all the line fluxes measured by pahfit.
+These fluxes have been corrected for dust attenuation by , where is the attenuation fraction from the pahfit model (dashed gray line in right panel of Figure 2).
+For the lines used in the upcoming sections, the attenuation-corrected fluxes are increased almost homogeneously by a factor of , meaning that the line ratios between these lines are almost unaffected by the dust attenuation assumed.
+It is worth noticing that the attenuation derived from pahfit of is mainly constrained by the two apparent absorption features at 9 and 18 , simultaneously.
+Fitting these wavelength ranges separately return a similar value of the attenuation, which excludes the presence of additional emission features, such as the 14 alumina oxidates reported in e.g.
+Hunt et al.
+(2025b).
 From the integrated spectrum shown in Figure 2, there is no clear evidence of any other PAH emission in CGCG 007-025.
 However, PAH features such as the 6.2, 7.7 or 8.6 are typically among the brightest ones (e.g., Hunt et al., 2010; Sandstrom et al., 2012; Chastenet et al., 2019).
-Following García-Bernete et al.
-(2022b), we obtain the upper limit on the 6.2 feature using the residuals from the pahfit model over the wavelength range 6.0 to 6.5 .
-We also report the upper limit on the 6.2 feature in Table 2.
-4.3 Ionized gas emission in the mid-IR
+In attempt to search for these other PAH features, we compare the integrated spectrum of CGCG 007-025 with the H II spectrum template from PDRs4All (Chown et al., 2024, see Figure 5 in the Appendix).
+While no clear evidence of emission is seen at 6.2 or 7.7 , the spectrum of CGCG 007-025 displays an excess of emission compatible with the additional presence of another PAH feature at 12.7 .
+We measure the flux of this feature by fitting a Drude profile to the continuum subtracted integrated spectrum.
+In order to obtain upper limits on the non-detected features, we follow the procedure described in Lai et al.
+(2025) and report the values in Table 1 (see Appendix B for more details).
+IV.3 Ionized gas emission in the mid-IR
 The integrated spectrum of the PAH-emitting region presented in Figure 2 reveals the detection of a plethora of nebular emission lines such as [Fe II] 5.34, [Ar II] 6.99, [Ne II] 12.81, [S III] 18.71, [Ar III] 8.99, [S IV] 10.51, [Ne III] 15.55, and [O IV] 25.89, as well as the high ionization line [Ne V]111Here we focus only on [Ne V] 14.32, since the accompanying line at 24.32 falls at the edge of ch4-medium and ch4-long, challenging the detection of the feature.
@@
 Accordingly, we search for the detection of other lines at the spaxel level, including lines of high ionization potential (IP) such as [O IV] and [Ne V].
-More concretely, following the methods outlined in Section 3.1, we focus on [Ne II], [Ne III], [O IV] and [Ne V] due to the diagnostic power of these lines (see Section 5).
-While [Ne II] and [Ne III] emission extend across the entire MIRI FoV (see inset in Figure 1), both [O IV] and [Ne V] emission are in a more confined configuration, co-spatial with the same compact region as the extended emission (see light green contour in Figure 2).
-4.4 Previous observations with Spitzer
-As part of an analysis on Blue Compact Dwarfs (BCDs) with Spitzer/IRS, Hunt et al.
-(2010) studied the infrared spectrum of CGCG 007-025.
-Combining the IRS low and high resolution gratings, they analyzed the integrated spectrum of the brightest star-forming region, and reported the detection of two PAH features: the 7.7 and 11.3 complexes.
-They also reported the detection of the fine structure lines [Ar III], [S IV], [Ne III] and [S III], together with some lines.
-We extracted the integrated MIRI/MRS spectrum of CGCG 007-025 using the methodology outlined in Section 3.2, with the spatial footprint of MIRI/MRS channel 1 (FoV 32 × 37) as the defining aperture to try to match the Spitzer/IRS aperture (see Figure 1).
-The top panel of Figure 3 displays this spectrum, with the different emission lines identified.
-In order to visually compare it with the literature, we degraded the MIRI/MRS data to match the spectral resolution of the IRS spectrum from Hunt et al.
-(2010).
-We convolved the MIRI/MRS integrated spectrum with the Spitzer/IRS LSF, and then binned the spectrum to a 0.03 wavelength step.
-This spectrum is shown as a maroon thick-line in both panels of Figure 3.
-The faint emission lines in the integrated MIRI/MRS spectrum (such as [Ar II] or [Ne V]) are not visible at the Spitzer sensitivity.
-Moreover, lines that are close to each other are blended at the Spitzer resolution.
-This is the case of the Pfund () and Humphrey () lines, which conspire to show up as a small bump at the position of the 7.7 PAH feature, clearly absent in MIRI/MRS high-resolution spectrum.
-This highlights the superior capability of MIRI/MRS compared to Spitzer to resolve mid-IR individual features.
-In the bottom panel of Figure 3 we show together the degraded MIRI/MRS spectrum with the reference spectrum from Hunt et al.
-(2010).
-We note a discrepancy in the spectral slopes, which may be due to a slight aperture mismatch and/or the specific sensitivity characteristics of MIRI.
-We also emphasized in green the theoretical central wavelengths of the PAH features.
-The two PAH complexes reported in Hunt et al.
-(2010) are marked in dark green.
-Based on this visual comparison and given the high spectral resolution and sensitivity of MIRI/MRS, we argue that only the 11.3 m feature is present in the galaxy, while the 7.7 m feature is not detected within the covered FoV.
-5 Discussion
-5.1 PAHs as star formation rate indicators
-PAH emission has long been proposed to be a potential star formation rate (SFR) indicator.
-Using Spitzer data, Xie & Ho (2019) presented a relation between the PAH luminosity and the SFR obtained from [Ne II] and [Ne III] for different types of galaxies, including BCDs, star-forming galaxies (SFGs) and AGN (active galactic nuclei).
-Although the PAH intensity is suppressed in the low-mass () galaxies, the Xie & Ho (2019) SFR-PAH calibrations can still be applied to this mass regime.
-Our MIRI/MRS data allows us to test for the validity of PAH features as SFR tracers, not only at low metallicity but also in a spatially resolved way.
-In order to estimate the SFR in the PAH emitting region, we used the prescription presented in Zhuang et al.
-(2019), which is an update of Ho & Keto (2007) to include the effects of metallicity.
-This estimator relies on the [Ne II] and [Ne III] luminosities and their fractional abundances, and includes a correction to account for the presence of [Ne V].
-The use of [Ne II] and [Ne III] to calculate the SFR is motivated by the lack of a reliable estimator based solely on mid-IR hydrogen lines.
-Moreover, this SFR tracer has been calibrated against the Paschen- emission, making it more suitable for the case explored here.
-Using the extinction corrected values from pahfit and a metallicity of (Berg et al., 2022; Fernández et al., 2023), we obtain a in the PAH emitting region.
-When compared with the literature, the point in CGCG 007-025 falls within the scatter of the other BCDs from Xie & Ho (2019) (see Figure 6 in Appendix B), and within the relation between for BCDs.
-However, this measurement only encompasses the region in which the PAH emission is reliably measured while, intriguingly, the area enclosing the SFR peak (solid white contour in Figure 2) does not display significant PAH emission.
-It remains possible that these calibrations do not hold in low-metallicity environments when using spatially resolved data.
-5.2 What PAH and line ratios reveal about CGCG 007-025
+More concretely, following the methods outlined in Section III.1, we focus on [Ne II], [Ne III], [O IV] and [Ne V] due to the diagnostic power of these lines (see Section V).
+While [Ne II] and [Ne III] emission extend across the entire MIRI FoV (see inset in Figure 1), both [O IV] and [Ne V] emission are in a more confined configuration, co-spatial with the same compact region as the extended emission (see light green contour and flux distribution in green in Figure 2).
+V Discussion
+V.1 What PAH and line ratios reveal about CGCG 007-025
 PAH features at different wavelengths can be used to trace the predominant ionization state of the overall PAH population.
@@
 Consequently, the ratios between PAH features provide insight into the dominant PAH population and the ionizing mechanisms within a galaxy.
-For instance, several studies have reported thresholds in the 7.7/11.3 and 6.2/11.3 ratios, with SF-dominated regions typically showing 6.2/11.3 larger than 1, while environments exposed to harder radiation fields (e.g., AGN) display ratios below 0.3 (e.g., Ogle et al., 2010; García-Bernete et al., 2022a; Donnan et al., 2023).
+For instance, several studies have reported thresholds in the 7.7/11.3 and 6.2/11.3 ratios222In this manuscript, we avoid the use of the 7.7 feature due to the large variations in the continuum noise around this feature (see Figure 5)., with SF-dominated regions typically showing 6.2/11.3 larger than 1, while environments exposed to harder radiation fields (e.g., AGN) display ratios below 0.3 (e.g., Smith et al., 2007; Ogle et al., 2010; García-Bernete et al., 2022b; Donnan et al., 2023).
 We show in Figure 3 the 6.2/11.3 PAH ratio against [Ne III]/[Ne II], which is tracing the level of ionization.
@@
 Seyfert galaxies move towards the center of the plot, showing an increase in the level of ionization and a further decrease in the 6.2/11.3 ratio.
-BCDs, on the other hand, cover a wider PAH ratio range while showing even higher ionization.
+Blue Compact Dwarfs (BCDs), on the other hand, cover a wider PAH ratio range while showing even higher ionization.
 CGCG 007-025 (orange arrow in the lower right corner) stands out as an outlier in this diagram, showing one of the largest [Ne III]/[Ne II] ratio together with a deficit on the 6.2/11.3 PAH ratio.
@@
 The location of CGCG 007-025 in the diagram may indicate the presence of a highly ionizing source, as its PAH strengths are consistent with those typically observed in AGN-dominated systems.
-Indeed, its PAH and mid-IR line properties closely resemble those observed in the nucleus of the Seyfert 2 galaxy NGC 7319 and in radio galaxies with strong jets, which show only a marginal detection of the 11.3 m PAH feature (Pereira-Santaella et al., 2022; García-Bernete et al., 2022b; Zakamska et al., 2016).
+Indeed, its PAH and mid-IR line properties closely resemble those observed in the nucleus of the Seyfert 2 galaxy NGC 7319 and in radio galaxies with strong jets, which show only a marginal detection of the 11.3 m PAH feature (Pereira-Santaella et al., 2022; García-Bernete et al., 2022a; Zakamska et al., 2016).
 Similarly, the [Ne III]/[Ne II] and PAH 6.2/11.3 ratios of the AGN NGC 5728 (García-Bernete et al., 2024) are consistent with those measured in this galaxy.
@@
 50 eV) and [Ne V] (I.P.
-97 eV), which are hard to explain in pure star-forming environments (e.g., Lutz et al., 1998; Schaerer & Stasińska, 1999).
+97 eV), which are hard to explain in pure star-forming environments (e.g., Lutz et al., 1998; Schaerer and Stasińska, 1999).
 Nevertheless, [O IV] emission is commonly observed in BCDs (e.g., Dale et al., 2009; Bernard-Salas et al., 2009), and recent MIRI/MRS observations have revealed extended [Ne V] in some extreme low-metallicity systems, such as SBS 0335–052E and I Zw 18 (Mingozzi et al.
 2025; Hunt et al.
-2025; see also e.g., Izotov et al.
+2025a; see also e.g., Izotov et al.
 2012, 2021 for detections on the optical).
@@
 Overlaid in green are the single stellar population (SSP) models from Martínez-Paredes et al.
-(2023) for an age222This is the only age grid that appears within the displayed [Ne V]/[Ne II], [Ne III]/[Ne II] and [O IV]/[Ne III] ranges and is the closest to the estimated age for CGCG 007-025 of 5 Myr from del Valle-Espinosa et al.
+(2023) for an age333This is the only age grid that appears within the displayed [Ne V]/[Ne II], [Ne III]/[Ne II] and [O IV]/[Ne III] ranges and is the closest to the estimated age for CGCG 007-025 of 5 Myr from del Valle-Espinosa et al.
 (2023) of 2.5 Myr, with a top-heavy IMF and and metallicity ranging from .
 Interestingly, in both panels, the data points are far away from the locus of the pure SF models.
-For the other two recently discovered [Ne V]-emitters - SBS 0335–052E (light blue markers and contours in Figure 5) and I Zw 18 - the only models capable of reproducing the observed line ratios require the presence of an additional source, such as an intermediate-mass black hole (IMBH; Mingozzi et al., 2025; Hunt et al., 2025).
+For the other two recently discovered [Ne V]-emitters - SBS 0335–052E (light blue markers and contours in Figure 4) and I Zw 18 - the only models capable of reproducing the observed line ratios require the presence of an additional source, such as an intermediate-mass black hole (IMBH; Mingozzi et al., 2025; Hunt et al., 2025a).
 Thus, we included in Figure 4 the IMBH grids from Richardson et al.
-(2022) with and the youngest age available (20 Myr), color-coded by the AGN fraction.
+(2025) with , an age of 5 Myr, limited to a metallicity range of , and color-coded by the AGN fraction.
 Consistently with Mingozzi et al.
@@
 However, these models fail to reproduce, simultaneously, the high and low ionization species seen in CGCG 007-025.
+We note, however, that exotic line ratios are not enough to probe the existence of an IMBH in this system, and additional observables should be tested.
+For example, the presence of broad components in the permitted lines, not detected in other forbidden lines, or the presence of X-ray emission, are characteristic of AGNs.
+In this vein, Senchyna et al.
+(2020) reported a prominent X-ray source located within the brightest SF region.
+Later, Senchyna et al.
+(2022) demonstrated that photoionization by extremely metal-poor stars, coupled with high-mass X-ray binaries or ultra luminous X-ray sources (ULXs), cannot reproduce the observed He II/H in this source.
+Moreover, in del Valle-Espinosa et al.
+(2023) we measured a broad component in the H line profile with velocity dispersion typical of AGN broad-line regions (1000 km/s), although the lack of other evidences of AGN activity at that time precluded the determination of its origin.
 Overall, we would like to remark this is the first time that the emission line ratios of this galaxy fall outside the SF locus in diagnostic diagrams (using e.g., UV and optical line ratios; Mingozzi et al., 2022; del Valle-Espinosa et al., 2023).
@@
 Stronger radiation fields would further destroy the PAHs, reducing the ratio below 5.
-For CGCG 007-025, we report a lower limit of , which would fall in the AGN demarcation.
-Nevertheless, this trend has been established in (super-)solar metallicity environments as well as in confirmed AGN host galaxies.At lower metallicities, it is still to be determined whether this ratio can be attributed to the harder ionizing spectrum (which would destroy the PAH complexes and increase the S(1) flux, simultaneously), or whether they are shaped simply because of the lack of PAH molecules in the first place.
+For CGCG 007-025, we report a upper limit of , which would fall in the AGN demarcation.
+Nevertheless, this trend has been established in (super-)solar metallicity environments as well as in confirmed AGN host galaxies.
+At lower metallicities, it is still to be determined whether this ratio can be attributed to the harder ionizing spectrum (which would destroy the PAH complexes and increase the S(1) flux, simultaneously), or whether they are shaped simply because of the lack of PAH molecules in the first place.
 It is worth noting that the PAH emission peaks at the location of high-IP lines ([O IV] and [Ne V]), with no significant PAH detections in regions outside of these high-IP boundaries.
@@
 For example, the formation of PAH has previously been seen in the dusty winds of massive and/or Wolf-Rayet (WR) stars (Lau et al., 2022; Taniguchi et al., 2025).
-Although WR stars seem capable of producing [Ne V] emission (see Tarantino et al., 2024; Hunt et al., 2025), current photoionization models of massive stars cannot reproduce the mid-IR line ratios seen in CGCG 007-025.
+Although WR stars seem capable of producing [Ne V] emission (see Tarantino et al., 2024; Hunt et al., 2025a), current photoionization models of massive stars cannot reproduce the mid-IR line ratios seen in CGCG 007-025.
 Further theoretical and observational efforts are thus required to understand the production of high-ionization photons and the formation of dust and complex molecules in low-metallicity starburst environments.
-6 Conclusions
+VI Conclusions
 In this letter, we present the first clear and spatially resolved PAH detection at in the dwarf galaxy CGCG 007-025 using the unique capabilities of JWST MIRI/MRS. Our findings are summarized below:
@@
 We explore the extension and location of this feature.
-The emission is spatially confined to a region of pc in the vicinity of the brightest clumps on the galaxy.
+The emission is spatially confined to a region of pc, offset from the brightest clumps on the galaxy.
 Using pahfit, we characterize the continuum and spectral features (nebular lines and PAHs) of the integrated spectrum of the PAH emitting region (right panel of Figure 2).
 We confirm the presence of the 11.3 PAH feature with a flux of .
-We do not find evidence of any other PAH feature in the integrated spectrum of CGCG 007-025. We place an upper limit detection on the 6.2 PAH complex of .
-The exclusive detection of the 11.3 feature indicates a PAH population dominated by larger, neutral molecules.
+We find evidence of another PAH feature at 12.7 , but the usually bright features at 6.2, 7.7 or 8.6 are absent in CGCG 007-025.
+The detection of the 11.3 and 12.7 feature may indicate a PAH population dominated by larger, neutral molecules.
 Moreover, the 6.2/11.3 ratio, in conjunction with a large [Ne III]/[Ne II] ratio (Figure 3), hint to the presence of a strong ionizing source, with [Ne III]/[Ne II] ratios compatible with those seen in harsh environments like AGNs or other BCDs.
@@
 At the time of reducing the data, the developer version of the pipeline was the only release which included the straylight correction step, a crucial step for correcting these intense cosmic ray events.
-Background subtraction is performed in Stage 3 using a master sky background created with the dedicated sky observations of each object.
+Background subtraction is performed in Stage 3 using a master sky background created with the dedicated sky observations.
 We opted for this option rather than a pixel-by-pixel correction, since the later introduces more noise in the final data products.
@@
 Besides our best efforts to correct from fringing and PSF undersampling, some spaxels situated at the peak of the mid-IR emission are still affected by these wiggling patterns.
-Appendix B Additional material
-In Table 1 we compile the complete list of emission lines included in pahfit for the modeling of the PAH emitting region.
-Table 2 displays the unattenuated emission line measurements from pahfit together with the flux of the 11.3 m PAH feature and the upper limit on the 6.2 m PAH feature.
-| Fine structure lines | H I lines | lines | ||||
-| Line name | IP | Transition | Line name | |||
-| [eV] | [] | [] | [] | |||
-| [Fe II] | 7.9 | 5.34 | Hu(10-6) | 5.129 | S(8) | 5.053 |
-| [Ar II] | 15.76 | 6.99 | Hu(9-6) | 5.908 | S(7) | 5.511 |
-| [Ar III] | 27.63 | 8.99 | Pf(6-5) | 7.45985 | S(6) | 6.109 |
-| [S IV] | 34.83 | 10.51 | Hu(8-6) | 7.503 | S(5) | 6.909 |
-| [Ne II]⋆ | 21.56 | 12.81 | 10-7 | 8.76 | S(4) | 8.026 |
-| [Ne V]⋆ | 97.12 | 14.32 | 9-7 | 11.3088 | S(3) | 9.665 |
-| [Ne III]⋆ | 40.96 | 15.55 | Hu(7-6) | 12.37188 | S(2) | 12.278 |
-| [S III] | 23.34 | 18.71 | 8-7 | 19.062 | S(1) | 17.035 |
-| [O IV]⋆ | 54.95 | 25.89 | ||||
-| [Fe II] | 7.9 | 25.99 | ||||
-| Fine structure lines | lines | PAH | |||
-| Line name | Flux | Line name | Flux | Feature | Flux |
-| [Fe II]5.34 | 0.4 0.3 | S(1) | 0.80 0.03 | 6.2 | |
-| [Ar II]6.99 | 0.29 0.08 | 11.3 | 5.1 0.4 | ||
-| [Ar III]8.99 | 3.86 0.08 | ||||
-| [S IV]10.51 | 31.2 0.4 | ||||
-| [Ne II]12.81 | 2.22 0.03 | ||||
-| [Ne V]14.32 | 0.38 0.02 | ||||
-| [Ne III]15.55 | 35.3 0.4 | ||||
-| [S III]18.71 | 43.4 0.4 | ||||
-| [O IV]25.89 | 24.4 0.4 | ||||
-| [Fe II]25.99 | 1.2 0.3 | ||||
-In Figure 6 we show the SFR-PAH relation for the 11.3 (left) and 6.2 (right) PAH features.
-The BCDs sample from Xie & Ho (2019) is displayed as purple diamonds, with open symbols representing upper limits on the PAH luminosities.
-The different SFR-PAH calibrations from Xie & Ho (2019) are shown in purple, with other calibrations from the literature displayed in green (Mordini et al., 2021).
-The measurements from CGCG 007-025 are shown in orange.
+We also verified that the astrometry of the MIRI/MRS datacubes matched the World Coordinate System (WCS) of the available HST data of this galaxy, as well as the WCS of the MUSE dataset from which we obtained the H contours.
+Appendix B Absence of typical PAH features in CGCG 007-025
+Despite of the bump at 11.3 m, the integrated spectrum of CGCG 007-025 does not display clear evidence of other PAH features.
+In Figure 5, we compare the integrated spectrum of CGCG 007-025 (gray line) with the PDRs4All template for HII regions (pink line, Chown et al., 2024).
+Both spectra have been continuum subtracted, and the template is also re-scale to match the flux of the 11.3 m feature.
+The spectrum of CGCG 007-025 shows no excess of emission at 6.2 and 7.7 m (left panel), indicating no evidence of these features in the aperture.
+However, the spectrum displays a slight excess of emission at 12.7 m, matching the aforementioned template (right panel).
+We measure the flux associated with the 12.7 feature on the continuum subtracted spectrum using a Drude profile.
+The integrated flux, corrected for attenuation, can be found in Table 1.
+Following Lai et al.
+(2025), we derive upper limits on the undetected PAHs by estimating the integrated fluxes of the 6.2, 7.7 and 8.6 bands.
+We define as the product of the width of the PAH profile and rms of the residuals from PAHFIT in the wavelength range 9.68–10.43 , where no PAH or emission lines are observed.
+The width of the PAH bands are defined as the interval spanning one FWHM, centered on the central wavelength (Smith et al., 2007).
+For the PAH complexes 6.2 and 7.7, the wavelength range expands from half a FWHM below the bluest subcomponent to half a FWHM above the reddest subcomponent.
+The upper limits on the PAH bands can be found on Table 1.
+Appendix C Additional material
+In Table 1 we compile the well-detected list of emission lines included in pahfit for the modeling of the PAH emitting region, as well as the unattenuated emission line measurements from pahfit.
+We also included the flux of the 11.3 and 12.7 m PAH features and the upper limits on the 6.2, 7.7 and 8.6 m PAH features.
+| Fine structure lines | lines | PAH | ||||||
+| Line name | IP | Flux | Line name | Flux | Feature | Flux | ||
+| [eV] | [] | [] | [] | [] | [] | |||
+| [Fe II] | 7.9 | 5.34 | 0.4 0.3 | S(1) | 17.035 | 0.80 0.03 | 6.2 | |
+| (0.5 0.3) | (1.00 0.03) | () | ||||||
+| [Ar II] | 15.76 | 6.99 | 0.29 0.08 | 7.7 | ||||
+| (0.3 0.1) | () | |||||||
+| [Ar III] | 27.63 | 8.99 | 3.86 0.08 | 8.6 | ||||
+| (5.8 0.1) | () | |||||||
+| [S IV] | 34.83 | 10.51 | 31.2 0.4 | 11.3 | 3.9 0.4 | |||
+| (49.0 0.6) | (5.1 0.4) | |||||||
+| [Ne II]⋆ | 21.56 | 12.81 | 2.22 0.03 | 12.7 | 7.0 0.4 | |||
+| (2.38 0.03) | (7.5 0.4) | |||||||
+| [Ne V]⋆ | 97.12 | 14.32 | 0.38 0.02 | |||||
+| (0.41 0.02) | ||||||||
+| [Ne III]⋆ | 40.96 | 15.55 | 35.3 0.4 | |||||
+| (40.3 0.5) | ||||||||
+| [S III] | 23.34 | 18.71 | 43.4 0.4 | |||||
+| (55.1 0.5) | ||||||||
+| [O IV]⋆ | 54.95 | 25.89 | 24.4 0.4 | |||||
+| (25.4 0.4) | ||||||||
+| [Fe II] | 7.9 | 25.99 | 1.2 0.3 | |||||
+| (1.2 0.3) | ||||||||